Please Help! Story Problems
1. The length of a walkway is 3 more than 5 times the width of the walkway. If the perimeter is 90 feet, find the dimensions of the walkway.
2. The width of a rectangular is 10 feet less than the length, find the length of each side if the perimeter s 100 feet.
3. Dan and Charlie have ban accounts with $96 between then. Dan has $2 less than Charlie in his bank accounts. How much does eac have in their bank accounts?
4. A pole 24 inches long was cut into 3 pieces. Max, Jeremy and Sherri each have a piece. Max had the shortest piece, Jeremy had a piece 2 inches longer, and Sherri had a piece 5 inches longer than Jeremy. How long was the pole each person ended up with.
Please Help

Answer by xcentaur(357)

1. The length of a walkway is 3 more than 5 times the width of the walkway. If the perimeter is 90 feet, find the dimensions of the walkway.

let the width be x feet.Then we understand that
length=(3+5x)feet
Perimeter=2(l+b)=2(x+3+5x)=2(6x+3)feet
then 90 feet=2(6x+3)
45=6x+3
6x=42
x=7 feet

width=7 feet,length=3+5(7)=38 feet.


2. The width of a rectangular is 10 feet less than the length, find the length of each side if the perimeter s 100 feet.

let the length be x feet.
then width=(x-10)feet
perimeter=100 feet=2(x+x-10)
100=2(2x-10)
50=2x-10
2x=60
x=30 feet

length=30 feet,width=20 feet


3. Dan and Charlie have bank accounts with $96 between then. Dan has $2 less than Charlie in his bank accounts. How much does eac have in their bank accounts?

let amount of charlie be x dollars.
then dan has = (x-2)dollars
given x+(x-2)=96
2x-2=96
2x=98
x=49

charlie has 49 dollars,dan has 47 dollars.

4. A pole 24 inches long was cut into 3 pieces. Max, Jeremy and Sherri each have a piece. Max had the shortest piece, Jeremy had a piece 2 inches longer, and Sherri had a piece 5 inches longer than Jeremy. How long was the pole each person ended up with.


let max's piece be x inches
then jeremy has (x+2) inches
sherri has (x+2)+5 inches
then x+(x+2)+(x+2)+5=24 inches
3x+9=24
x+3=8
x=5 inches


Max=5 inches,Jeremy=7 inches,Sherri=12 inches


Good luck,
Prabhat.